What is the outside surface area of a solid cylinder 8 inches
high with a diameter of 4 inches?

Answer by stanbon(75887)

What is the outside surface area of a solid cylinder 8 inches
high with a diameter of 4 inches?
--------------------
Surface area = (top + bottom + side)
top = (pi)2^2= 4pi
bottom = (pi)2^2 = 4pi
side = (pi)4*8 = 32pi
----------------------------
Surface area =- 40pi sq inches
================================
Cheers,
Stan H.